A = P 8500(1 + .09)-12
P = 8500(1.09)-12
P = 8500/(1.09)-12
I am trying to determine how much principal I need to invest today, to reach $8500 in 12years at 9% compounded interests. I can't seem to figure out the principal that I need - thank you!

Answer by mananth(16946)

Principal P = x
Amount= 8500
years=n 12
compounded 1 times a year
Rate = 9.00 0.090
Amount = P*((n+r)/n)^n

8500 = x *( 1 + r/n )^ n


8500 = P ( 1 + 0.090 / 1 ) ^ 1 * 12

8500 = P( 1.09000 )^ 12

8500 /( 1.09 )^ 12 = P

P= $3022.05